OpenStudy (anonymous):

I really need help :/ I will fan and give medal It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

OpenStudy (misty1212):

Hi!

OpenStudy (misty1212):

we can do this but it will take a minute or two ready?

OpenStudy (misty1212):

this is where you say "yes, i am ready" or "no forget it"

OpenStudy (anonymous):

Yes I am sorry computer glitched

OpenStudy (misty1212):

oh ok i didn't do it in advance but we can work it out i am sure

OpenStudy (misty1212):

lets say it takes john \(x\) hours so it takes brian \(x+15\) hours that means john's rate is \(\frac{1}{x}\) per hour and brians is \(\frac{1}{x+15}\) per hour now we have to find the combined rate, which we do by adding

OpenStudy (misty1212):

the combined rate is \[\frac{1}{x}+\frac{1}{x+15}=\frac{2x+15}{x(x+15)}\] and it take them 4 hours that rate to do one job, so we have to solve \[\frac{2x+15}{x(x+15)\times 4=1\]

OpenStudy (misty1212):

oops \[\frac{2x+15}{x(x+15)}\times 4=1\]

OpenStudy (misty1212):

let me know where i lost you solving this equation is not so hard \[4(2x+15)=x(x+15)\\ 8x+60=x^2+15x\] and then you have a quadratic equation to solve

OpenStudy (anonymous):

Oh okay I see now. Someone else got the same quadratic equation I just wasn't sure how they got it until now

OpenStudy (misty1212):

ok then i guess it is right!

OpenStudy (anonymous):

Yes and the number of hours it takes Brian in 20 hours

OpenStudy (anonymous):

is*

OpenStudy (misty1212):

\[x^2+7x-60=0\\ (x-5)(x+12)=0\\ x=5\]

OpenStudy (misty1212):

so john took 5 hours, brian 20

OpenStudy (anonymous):

Thank you so much for your help :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]